### v1

# Scaling Properties of Speech Language Models

Abstract
Speech Language Models (SLMs) aim to learn language from raw audio, without textual resources. Despite significant advances, our current models exhibit weak syntax and semantic abilities. However, if the scaling properties of neural language models hold for the speech modality, these abilities will improve as the amount of compute used for training increases. In this paper, we use models of this scaling behavior to estimate the scale at which our current methods will yield a SLM with the English proficiency of text-based Large Language Models (LLMs). We establish a strong correlation between pre-training loss and downstream syntactic and semantic performance in SLMs and LLMs, which results in predictable scaling of linguistic performance. We show that the linguistic performance of SLMs scales up to three orders of magnitude more slowly than that of text-based LLMs. Additionally, we study the benefits of synthetic data designed to boost semantic understanding and the effects of coarser speech tokenization.
1 Introduction
Figure 1: Speech Language Models test loss curves for all our single-epoch runs. Axes are in logarithmic scale. The envelope of minimal loss per FLOP (black dots) follows a power law (dashed line).
Inspired by the remarkable ability of preschool children to learn language from raw sensory inputs, Lakhotia et al. (2021) introduced in their seminal paper the textless NLP (Natural Language Processing) project. The project aimed to leverage advances in self-supervised speech representation learning for unsupervised unit discovery Hsu et al. (2021); Chung et al. (2021) and generative neural language models Brown et al. (2020) to jointly learn the acoustic and linguistic characteristics of a language from audio alone, without access to textual supervision (e.g. lexicon or transcriptions). They formalized this goal in the task of Generative Spoken Language Modeling (GSLM), in which a language model is trained on sequences of self-supervised learned speech units.
Despite a significant body of research on these speech-based language models (SLMs) Lakhotia et al. (2021); Kharitonov et al. (2022); Borsos et al. (2023); Hassid et al. (2023), they are still far from matching the syntactic and semantic abilities of text-based systems Hassid et al. (2023). Therefore, the promise of textless NLP is yet to be realized. However, if the scaling behavior of text-based neural language models Brown et al. (2020); Kaplan et al. (2020) holds for the speech modality, we can reasonably expect those abilities to improve as the amount of compute used for training increases.
In this work, we apply recently proposed models of the scaling behavior of neural language models to SLMs, and use them to estimate the scale at which our current methods will match the linguistic performance of Large Language Models (LLMs), generative text-based systems that have achieved remarkably strong performance across a wide range of NLP applications Brown et al. (2020). The main contributions of this work are:
Figure 2: Downstream linguistic performance scaling with compute for LLMs and SLMs. Axes are in logarithmic scale. Syntactic (BLIMP) and semantic (Topic Cloze and Story Cloze) metrics follow a power law before starting to saturate. Linguistic performance scales up to three orders of magnitude more slowly in SLMs relative to LLMs.
  • We trained over 50 SLMs with different number of parameters and data budgets. We show that the test loss of SLMs follows scaling power laws as those observed in text-based LLMs (Figure 1), and use the methods from Hoffmann et al. (2022) and Muennighoff et al. (2023) to model the scaling behavior of SLMs.
  • We establish a strong correlation between the test loss of neural LMs and the downstream metrics commonly used to evaluate their syntactic and semantic abilities. Therefore, the linguistic performance of LMs follows similar scaling laws (Figure 2). We leverage this insight to determine the relative efficiency with scale of SLMs relative to LLMs.
  • We speculate that SLMs require more context than fits in their context window to acquire from commonly used speech datasets the semantic understanding measured by our metrics. Accordingly, we propose a new speech dataset to boost semantic understanding in SLMs. Specifically, we synthesized a spoken version of the Tiny Stories dataset Eldan and Li (2023), and show that its use during pre-training improves semantic downstream performance.
  • On the basis of our previous observation, we studied the use of unigram tokenization to shorten sequences and pack more information in the context window of SLMs. However, our results suggest that a coarser tokenization is detrimental to downstream performance.
2 Background
2.1 Generative spoken language modeling
We follow the GSLM framework from Lakhotia et al. (2021). The general GSLM pipeline is composed of three separately trained models: (i) a speech tokenizer, (ii) a language model, and (iii) a vocoder (token-to-waveform) module. In the following, we provide background for the speech tokenizer and LM, as these are the components we use in this work. For details about the vocoder please refer to Lakhotia et al. (2021).
Speech tokenizers transform raw speech waveforms into discrete representations. A speech encoder is used to extract continuous representations that are then transformed into discrete sequences through vector quantization. Formally, let $\mathcal{X}\in\mathbb{R}$ denote the domain of audio samples, a waveform is therefore a sequence of samples $x=(x_{1},\dots,x_{T})$, where $x_{t}\in\mathcal{X}$ for all $1\leq t\leq T$. An encoder $F:\mathcal{X}^{m}\to\mathbb{R}^{d}$ transforms windows of samples of width $m$ into $d$ dimensional continuous frame representations. Applying $F$ to $x$ yields a sequence of frame representations $z=(z_{1},\dots,z_{T^{\prime}})$, where usually $T^{\prime}<T$. Subsequently, a k-means algorithm is applied to the encoder output to generate a sequence of discrete speech tokens $u=(u_{1},\dots,u_{T^{\prime}})$, where $u_{i}\in\{1,\dots,K\}$ for $1\leq i\leq T^{\prime}$, and $K$ is the size of the vocabulary.
Language models aim to learn the joint probability of token sequences $P(w_{1},\dots,w_{n})$. By the chain rule of probability, the probability of a sequence can be computed as a product of its conditional probabilities:
$P(w_{1},\ldots,w_{n})=\prod_{i=1}^{n}P(w_{i}|w_{1},\ldots,w_{i-1})$ (1)
Neural LMs, parameterized by $\theta$, are neural networks that model the conditional probabilities $P_{\theta}(w_{i}|M(w_{1},\ldots,w_{i-1}))$, where $M$ is a representation of the previous tokens. The network is optimized to minimize the negative log-likelihood of observed ground truth sequences:
$L=-\sum_{i=1}^{n}P_{\theta}(w_{i}|M(w_{1},\ldots,w_{i-1}))$ (2)
Nowadays, the network is typically a transformer Vaswani et al. (2017). LLMs are large transformer LMs trained on large text corpora (billions of parameters and tokens). SLMs are neural LMs applied to speech tokens $u$.
2.2 Scaling laws for neural language models
The performance of deep learning models often behaves predictably as a function of model size, dataset size, and compute Hestness et al. (2017). Kaplan et al. (2020) showed that the loss $L$ (Equation 2) of large neural LMs scales with a power law behavior as a function of these three scale factors:
$L(C)\propto C^{\gamma},\quad L(N)\propto N^{\alpha},\quad L(D)\propto D^{\beta}$ (3)
Where $C$ is the amount of compute (in FLOPS), $N$ is the number of parameters of the model, and $D$ is the number of training tokens.
Building upon their work, Hoffmann et al. (2022) proposed a parametric function to model the final loss of neural LMs trained for a single epoch as a function of $N$ and $D$:
$\hat{L}(N,D)=E+\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}$ (4)
Where the first term is the loss for an ideal LM, and should correspond to the entropy of the distribution of token sequences. The second term captures the approximation error that results from using a neural network with $N$ parameters to approximate the ideal generative process. The final term reflects that the model is not trained to convergence, as a finite number of optimization steps are performed on a sample of size $D$ from the real distribution.
Hoffmann et al. (2022) aimed to solve the problem of optimal allocation of resources given a fixed compute budget $C_{\text{avail}}$. They proposed to approximate the compute needed to train a transformer LM with $N$ parameters on $D$ tokens as $C\approx 6ND$. Then, the problem of optimal allocation of compute for model size and training data is:
$\min_{N,D}\hat{L}(N,D),\quad\text{s.t.}\quad 6ND=C_{\text{avail}}$ (5)
For which the solution is:
$N_{\text{opt}}(C)$ $=G\left(\frac{C}{6}\right)^{a}$ (6)
$\,D_{\text{opt}}(C)$ $=\frac{1}{G}\left(\frac{C}{6}\right)^{b}$
With:
$G=\left(\frac{\alpha A}{\beta B}\right)^{\frac{1}{\alpha+\beta}},\,a=\frac{\beta}{\alpha+\beta},\text{ and }b=\frac{\alpha}{\alpha+\beta}$
Muennighoff et al. (2023) generalized Equation 4 to the case of multi-epoch training by replacing $D$ and $N$ with terms corresponding to the effective data $D^{\prime}$ and effective model parameters $N^{\prime}$:
$\hat{L}(N^{\prime},D^{\prime})=E+\frac{A}{N^{\prime\alpha}}+\frac{B}{D^{\prime\beta}}$ (7)
Where $D^{\prime}\leq D$ is the number of effective training tokens, assuming that the value of repeated tokens decays exponentially. Similarly, they note that oversized models offer diminishing returns per parameter, as excess parameters learn the same features and do not add value (in the extreme). They propose an exponential decay model for them, yielding a number of effective parameters $N^{\prime}\leq N$. They derived the expressions for $D^{\prime}$ and $N^{\prime}$ as:
$D^{\prime}$ $=U_{D}+U_{D}R_{D}^{*}(1-e^{\frac{-R_{D}}{R_{D}^{*}}})$ (8)
$N^{\prime}$ $=U_{N}+U_{N}R_{N}^{*}(1-e^{\frac{-R_{N}}{R_{N}^{*}}})$
Where $U_{D}$ is the number of unique tokens used, $R_{D}=\frac{D}{U_{D}}-1$ is the number of repetitions (0 for a single epoch), $U_{N}$ is the number of parameters needed to optimally fit $U_{D}$ according to Equation 6, $R_{N}=\frac{N}{U_{N}}-1$ is the number of excess parameters, and $R_{D}^{*}$ and $R_{N}^{*}$ are constants.
The constants $E$, $A$, $B$, $\alpha$, $\beta$, $R_{D}^{*}$ and $R_{N}^{*}$ can be estimated empirically by fitting Equation 4 or 7 to a set of tuples $(N,D,R_{N},R_{D},L)$ obtained from training experiments with different budgets.
3 Experimental setup
3.1 Models and training
We adhere to the framework described in Section 2.1. For the speech tokenizer, we use a pre-trained HuBERT model Hsu et al. (2021) with frame-rate of 25 Hz as the speech encoder $F$, and a vocabulary size of $K=500$. This setup reports the best performance among publicly available models Hassid et al. (2023). For the SLMs we use the Llama architecture Touvron et al. (2023) with context window of 2050 tokens. Table 1 describes the model sizes used in our experiments. For the LLMs, we use the Pythia suite of pre-trained LLMs Biderman et al. (2023), ranging in size from 14M to 6.9B parameters (we do not use the largest 12B model), and trained with $\sim$300B tokens.
All SLMs are optimized using AdamW Loshchilov and Hutter (2019) with weight decay of 0.1, maximum learning rate of 5e-4, half-cycle cosine decay learning rate schedule to 5e-5, and a warm-up initial stage of $\max(100,0.01\,n_{iters})$ steps, where $n_{iters}$ is the number of training steps, which varies for each experiment according to the data budget. We use batch sizes of 64, 128, 256 and 512 for the models with 20M, 85M, 155M and 309M, and 828M parameters, respectively.
Size Layers Model dim. Heads
20M 6 512 8
85M 12 768 12
155M 12 1024 16
309M 24 1024 16
823M 16 2048 32
Table 1: Models description.
To fit the constants in Equations 4 and 7, we adopt the approaches of Hoffmann et al. (2022) and Muennighoff et al. (2023), utilizing the Huber loss with $\delta=0.03$ as the error function and L-BFGS as optimizer. Following Muennighoff et al. (2023), we first fit the parameters $E$, $A$, $B$, $\alpha$, and $\beta$ using the single-epoch runs, and afterwards fit $R^{*}_{D}$ and $R^{*}_{N}$ using the multi-epoch runs.
3.2 Evaluation
We use the sBLIMP task Nguyen et al. (2020) to measure syntactic performance. In sBLIMP, the model is presented with a matched pair of speech segments, grammatical and ungrammatical sentences. The objective is to assign higher probability to the grammatical sentence.
To evaluate semantic understanding we use the spoken StoryCloze benchmark from Hassid et al. (2023), a spoken version of the StoryCloze textual benchmark Mostafazadeh et al. (2016), which consists of 4k five-sentence commonsense stories. In StoryCloze, the model receives as input the first four sentences of a story, and has to assign higher probability to the correct final sentence than to an adversarial negative sample. The spoken benchmark comes in two versions: Story Cloze and Topic Cloze. The difference between them lies in how the negative sample is generated. Spoken Story Cloze uses the same samples as the textual benchmark, which require commonsense reasoning to distinguish from the real ending. In Topic Cloze, the negatives are randomly sampled from the whole dataset, and therefore measures the ability of the model to stay on topic.
Regarding upstream performance, in all cases we report and use for the parametric fits the average loss (Equation 2) on the test set.
3.3 Data
Dataset Hours HuBERT Tokens Unigram
LibriSpeech 960 67M 38M
LibriLight 53k 3.74B 2.11B
SWC 1k 32M 19M
Tedlium 1.6k 0.11B 67M
People 7k 0.48B 0.29B
Vox Populi 24k 1.64B 1.08B
sTinyStories 72k 4.82B 2.71B
Total 160k 10.89B 6.31B
Table 2: Datasets statistics. The Unigram column corresponds to the dataset of HuBERT tokens compressed through unigram tokenization.
3.3.1 Datasets
We use a collection of publicly available English speech datasets for training: LibriSpeech Panayotov et al. (2015), LibriLight Kahn et al. (2020), SWC Baumann et al. (2019), Tedlium Hernandez et al. (2018), People’s Speech Galvez et al. (2021), and Vox Populi Wang et al. (2021b); and a novel dataset: sTinyStories, a spoken version of the Tiny Stories dataset Eldan and Li (2023) that we synthesized using the single-speaker TTS system provided by Wang et al. (2021a). Tiny Stories is a synthetic text corpus of short stories designed to boost commonsense reasoning in neural LMs. We propose sTinyStories because we hypothesize that the semantic understanding that tasks such as Story Cloze measure is hard to acquire from commonly used speech datasets. Consider for instance the audiobooks in LibriLight. The data has long-range dependencies spanning multiple pages, whereas our SLMs can ingest roughly a dozen sentences of spoken text in their context window. Other datasets, which were mainly designed to serve as training data for automatic speech recognition systems, consist of too small fragments of audio that lack meaningful causal structure. sTinyStories consists of full stories with causal structure that fit within the context window of our SLMs.
We do not include samples from sTinyStories in our test set, as we intend to use our test loss as measure of the quality with which SLMs model natural language, not synthetic one. For other datasets we use the defined held-out sets for testing. In cases where a held-out set is not defined, we randomly sampled 1% of the data to serve as test set. See Table 2 for dataset sizes.
3.3.2 Data budgets
In order to have a representative set of samples to fit Equations 4 and 7, for each model size, we performed training runs with a ratio of training tokens $D$ to parameters $N$: $D/N\in\{2,4,8,16,32,64,128\}$. This setup yields single-epoch and multi-epoch runs for the larger models (e.g. for the model with 155M parameters the maximum number of training tokens corresponds to 1.82 epochs), but not for the smaller models (e.g. for the model with 85M parameters the maximum number of training tokens corresponds to 0.99 epochs). To have better samples to fit Equation 7, we did additional experiments so that for each model size there were runs with training epochs in $\{2,4,8,10\}$.
4 Results
4.1 Gains from sTinyStories
Figure 3: Gains from synthetic data on downstream semantic performance of SLMs. Pre-training on sTinyStories yields consistent improvements on semantic understanding relative to pre-training on audiobooks (LibriSpeech plus LibriLight). Performance gains hold for mismatched train and test speakers.
In order to determine if sTinyStories meaningfully contributes to the semantic understanding of SLMs, we compare the performance on Topic Cloze and Story Cloze of models trained on one epoch of the union of LibriSpeech and LibriLight, against models trained on an equivalent amount of sTinyStories tokens. Figure 3 shows the obtained results. Models trained on sTinyStories consistently outperform those trained on audiobooks across all model scales. A factor that could contribute to the observed performance gain is the match between training and evaluation speakers, as both sTinyStories and Story Cloze were synthesized using the single-sepaker TTS from Wang et al. (2021a). However, we believe this to be unlikely as the speech tokenizer we use likely captures little speaker-specific information Nguyen et al. (2023). To isolate the potential impact of speaker mismatch between training and evaluation data, we created a multi-speaker version of the Story Cloze benchmark using Bark TTS  [^1] https://github.com/suno-ai/bark, and repeat the evaluations. The results, also shown in Figure 3, indicate that even with mismatched train and test speakers training on sTinyStories yields performance gains.
4.2 Scaling laws
We trained multiple SLMs for each model size with different data budgets as described in Section 3.3.2. The resulting learning curves for single-epoch runs are presented in Figure 1 as a function of compute, and show that the envelope of minimal loss per FLOP follows a power law.
4.2.1 Downstream scaling with compute
We analyze the relationship between the upstream and linguistic downstream performance in SLMs and LLMs. Figure 4 shows the obtained results. Syntactic and semantic downstream metrics before saturation are strongly correlated with the upstream test loss in both LLMs and SLMs. Therefore, the envelope of maximum downstream performance per FLOP also follows a power law, i.e. for a downstream performance function $Q$, $Q\propto C^{\gamma_{q}}$. The power laws for the different performance metrics are depicted in Figure 2 and the exponents presented in Table 3.
Figure 4: Correlation between downstream linguistic performance and test loss for LLMs and SLMs. Syntactic (BLIMP) and semantic (Topic Cloze and Story Cloze) metrics are strongly linearly correlated with the upstream test loss before saturation.
These results allow us to compare the efficiency with scale of LLMs and SLMs. For each metric, we can interpret the ratio between the $\gamma_{q}$ exponents of the power laws of LLMs and SLMs as the relative efficiency with scale. For BLIMP, the ratio is $\frac{0.066}{0.021}=3.14$, indicating that for an increase in compute $\Delta C$ yielding a $\Delta Q$ in LLM’s syntactic performance, SLMs require $10^{3.14}\Delta C$ to get the same $\Delta Q$. Similarly, for Topic Cloze and Story Cloze the ratios are $1.56$ and $2.7$, respectively.
Modality $\gamma_{q}$
BLIMP tCloze sCloze
Text 0.066 0.039 0.046
Speech 0.021 0.025 0.017
Table 3: $\gamma_{q}$ power law coefficients of downstream performance with compute as depicted in Figure 2.
4.2.2 Scaling with parameters and tokens
We fitted the functions from Equations 4 and 7 to our data using the procedure described in Section 3.1. We present the empirically fitted scaling law parameters and compare them to the ones obtained for text by Muennighoff et al. (2023) in Table 4. From Equation 6, $N_{opt}\propto C^{a}$ and $D_{opt}\propto C^{b}$. For both modalities $a\approx b\approx 0.5$, suggesting that as compute increases, model size and data should be scaled equally for optimal performance. Contrary to text, $R^{*}_{N}>R^{*}_{D}$, indicating that repeated tokens decay faster than excess parameters (albeit both slower than in text). Therefore, in SLMs, compute allocated to parameters should scale faster than compute allocated for epochs.
E A B $\alpha$ $\beta$ $R^{*}_{N}$ $R^{*}_{D}$
Text Muennighoff et al. 1.87 521 1488 0.35 0.35 5.31 15.4
Speech 1.73 13.9 39.8 0.25 0.24 31.0 25.0
Speech (Unigram) 1.42 3.85 8.90 0.15 0.16 - -
Table 4: Scaling law parameters fit to Equations 4 and 7 for different language tokenizations.
Figure 5: Comparison of the scaling behavior of SLMs trained on raw speech tokens and unigram compressed tokens. Axes are in logarithmic scale. The upstream loss of SLMs trained on unigram tokens scales better with compute, but downstream performance scales worse. Notably, the Story Cloze metric for SLMs trained on unigram tokens does not seem to improve with increased compute.
4.3 Unigram tokenization
As mentioned in Section 3.3, we believe that the limited context window of SLMs could cripple their ability to model the long-range dependencies in language required for causal reasoning. Seeking to mitigate this limitation, we apply unigram tokenization to shorten the length of speech token sequences. We use the SentencePiece tokenizer Kudo and Richardson (2018) with a vocabulary size of 5000. We choose the vocabulary size on the scale of previous works that have used similar tokenization strategies for speech applications Chang et al. (2023). The resulting dataset sizes after compression are presented in Table 2.
We train a set of Speech LMs on the compressed datasets, with model sizes up to 309M parameters and data budgets ranging from 740M to 6.31B tokens. We analyze the scaling behavior of the upstream and downstream metrics and compare it with SLMs trained on raw HuBERT speech tokens in Figure 5. SLMs trained on unigram compressed speech tokens show similar upstream scaling with compute, but worse downstream scaling. Notably, the performance on the StoryCloze benchmark does not seem to scale with compute.
We fitted the function from Equation 4 to the results obtained on the compressed dataset. Table 4 presents the resulting scaling law parameters. Similar to the previous findings, for a given compute budget, scaling model size and training data equally is optimal for performance. Due to the poor downstream results obtained with unigram tokenization and the lack of sufficient compute resources, we did not perform multi-epoch training experiments.
5 Related work
Previous works have studied the scaling behavior of neural networks on speech applications. Droppo and Elibol (2021) showed that acoustic models trained with an auto-predictive coding loss follow similar power laws to those observed in neural LMs. Aghajanyan et al. (2023) used the scaling laws from Hoffmann et al. (2022) to model the scaling behavior of the upstream loss of neural LMs on multiple modalities, including speech. They used a speech tokenizer with higher framerate (50 Hz) and vocabulary size ($K=2000$) than the one we used (Section 3.1). Such fine-grained tokenizers capture a lot of the paralinguistic information in speech Nguyen et al. (2023). Therefore, their speech tokens can be considered almost a different modality due to the acoustic variance. Furthermore, they do not study the behavior with scale of downstream performance. In this work, we focus on the linguistic content of the signal. As reported by Hassid et al. (2023), our speech tokenizer performs best on downstream linguistic applications, and is therefore a more suitable choice to study the scaling behavior of the linguistic performance of SLMs.
This paper is perhaps most closely related to the work of Hassid et al. (2023). We largely follow their setup in terms of model architecture and evaluation metrics. Similar to our work, they reported improved linguistic downstream performance with scale in SLMs. However, they did not characterize their scaling behavior. Our models of the scaling behavior of SLMs allow practitioners to determine the compute needed to attain a specific loss, syntactic and/or semantic performance; and the optimal allocation of that compute with respect to parameters and tokens. To the best of our knowledge, we are the first to model the scaling properties of downstream linguistic performance in SLMs, and to study the scaling of the considered downstream metrics on text-based LLMs. This enables a comparison between the two modalities in terms of scaling efficiency.
6 Discussion
6.1 On our results
Our work showed that the upstream and downstream linguistic performance of our current methods for GSLM scales predictably with compute. This suggests that with sufficient computational resources, the goal of the textless NLP project of achieving neural LMs trained exclusively on speech, and matching the linguistic proficiency of their text-based counterparts, is achievable. The cost of such models could be prohibitive though, as we estimate that they will require up to three orders of magnitude more compute than a text-based LLM to achieve equivalent performance. In this regard, recent methods that leverage transfer learning from text-based LLMs Hassid et al. (2023); Zhang et al. (2023); Nguyen et al. (2024) are likely to be a better choice to achieve highly performant generative speech models. These hybrid text-speech generative models often enable cross-modal applications such as ASR or TTS. However, it remains to be seen how knowledge transfer from LLMs performs when the speech data is in a different language than the one the LLM was trained on. If there is no significant cross-lingual knowledge transfer between text and speech modalities, SLMs could still be an attractive choice for low-resource languages.
We explored the use of synthetic data and coarser tokenization to increase the semantic abilities of SLMs. Our synthetic dataset improved semantic performance, but using a coarser tokenization led to overall degradation of downstream performance. We do not have yet an hypothesis for why coarser tokens degrade performance, as this seems counter-intuitive, and contradicts the findings on other speech applications Chang et al. (2023). We leave this as an interesting issue to address in future work. Moreover, we believe that working on methods that allow to increase the information density per context-window of SLMs holds promise to improve their scaling behavior.
6.2 On limitations
Any extrapolation from our models of the scaling behavior of SLMs should be considered optimistic for the following reasons:
  1. 1.
    The LLMs from the Pythia suite that we used in this study are likely overtrained (all models were trained with $\sim$300B tokens). Optimally trained LLMs (according to Equation 6) should show better performance with scale, and therefore widen the gap with the scaling efficiency of SLMs.
  2. 2.
    The envelope of minimal loss per FLOP (Figure 1) might show a slight negative curvature at larger scale Hoffmann et al. (2022), reducing the scaling efficiency.
  3. 3.
    Our models for downstream performance ignore the fact that the metrics saturate. As observed in text LLMs, the improvements with scale slow down as performance approaches the saturation value. It is likely that, due to saturation, the compute required to yield a particular performance will be larger than predicted. Moreover, due to the lower density of linguistic information per context window in SLMs relative to LLMs, the saturation values of the metrics may be lower for SLMs.
7 Conclusions
We have trained a large set of SLMs of different sizes and on different data budgets. Using the collected data from those experiments, we studied the scaling properties of their upstream and downstream performance using recently proposed models of scaling laws for neural LMs. The obtained models of the scaling behavior allow practioners to optimally allocate compute to attain a specific loss, syntactic and/or semantic performance. We showed that the pre-training loss and downstream linguistic performance of SLMs and LLMs is highly correlated, and that they both scale predictably according to power laws. This predictable behavior allowed us to compare the scaling properties of SLMs and LLMs, from which we established that the linguistic abilities of SLMs scale up to three orders of magnitude more slowly than those of LLMs. Additionally, we proposed a new speech dataset, sTinyStories, and showed that its use during pre-training improves downstream semantic performance in SLMs. Finally, we explored the use of coarser speech tokenizations as a method to increase the ability of SLMs to model long-range dependencies. However, our results suggest that this is detrimental to downstream performance.
Acknowledgements
We are grateful to the French National Research Agency for their support through the ANR-20-CE23-0012-01 (MIM) grant, and the Institute of Convergence ILCB, supported by grants from France 2030 (ANR-16-CONV-0002) and the Excellence Initiative of Aix-Marseille University (A*MIDEX). This work was granted access to the HPC resources of GENCI-IDRIS under the allocation AD011014044.